# Scattering resonances and pairing in a Rabi-coupled Fermi gas

### v1

@@ Abstract
$G_{\sigma\sigma^{\prime}}({\bf k},\omega)=\bra{0}\hat{f}_{{\bf k}\sigma^{\prime}}(\omega-\hat{H}_{0}+i0)^{-1}\hat{f}_{{\bf k}\sigma}^{\dagger}\ket{0},$ (5)
with $\sigma,\sigma^{\prime}=\{1,2\}$, are not diagonal and depend on the Rabi drive (the factor $+i0$ shifts poles infinitesimally into the lower half place). The diagrams corresponding to the Rabi-dressed Green’s functions are presented in Fig. 2(a,b). By contrast, the third species is unaffected by the Rabi drive and thus its single-particle Green’s function is simply given by $G_{3}({\bf k},\omega)=(\omega-\epsilon_{{\bf k}3}+i0)^{-1}$.
@@
We now turn to the interactions between an atom of species 3 with a Rabi-coupled atom, Fig. 1(a). Although the form of $\hat{V}$ does not include direct interactions between species 1 and 2, the Rabi-induced interconversion processes discussed above must be accounted for, since it can couple bound and scattering states. As a result, we find that the two-body $T$-matrix that describes the interactions can be conveniently written as a $2\times 2$ matrix equation of the form
@@
$\mathbf{g}$ $=\begin{pmatrix}g_{13}&0\\ 0&g_{23}\end{pmatrix},\leavevmode\nobreak\ \leavevmode\nobreak\ \mathbf{\Pi}(\omega)=\begin{pmatrix}\Pi_{11}(\omega)&\Pi_{12}(\omega)\\ \Pi_{21}(\omega)&\Pi_{22}(\omega)\end{pmatrix},$ (7)
@@
$\Pi_{\sigma\sigma^{\prime}}(\omega)$ $=\int\frac{d\omega^{\prime}}{2i\pi}\sum_{\bf q}G_{\sigma\sigma^{\prime}}(-{\bf q},-\omega^{\prime})G_{3}({\bf q},\omega+\omega^{\prime}).$ (8)
@@
Figure 3: Rabi-dressed Feshbach resonances and hybrid bound states as a function of detuning $\delta/\Omega$. (a) Dressed scattering length for three sets of bare scattering lengths ($a_{13},a_{23}$) corresponding to the points marked by (,$\bullet$,$\blacklozenge$,$\blacksquare$ , ◆ , ■) symbols in Fig. 1(b). (b,c) Energy spectrum for $(a_{13},a_{23})\sqrt{2m_{r}\Omega}=(-0.3,0.8)$ in (b), and $(a_{13},a_{23})\sqrt{2m_{r}\Omega}=(0.6,0.8)$ in (c). We show the hybrid bound states (black lines), with the dashed red and blue lines corresponding to the bare 1-3 and 2-3 bound states. The purple lines show the dressed single-particle energies ϵ±subscriptitalic-$\epsilon_{\pm}$ and the shaded area highlights the scattering continuum of the driven system.
@@
In particular, we see that in general $a_{-}\neq a_{13}c^{2}+a_{23}s^{2}$, and that the effective range $r_{\text{eff}}\leq 0$, similar to the case for a standard two-channel model of a magnetically tunable Feshbach resonance [40, 41]. We also note that when $a_{23}=0$, Eq. (10a) reduces to the effective scattering length derived in Refs. [42, 43]. In the high-symmetry scenario where $a_{13}=a_{23}=a$, Eq. (10a) reduces to $a_{-}=a$, and thus for this special case, the dressed scattering length is independent of the Rabi drive and only diverges when $a$ is itself divergent  [^1] Indeed, when $a_{13}=a_{23}=a$, the $T$ matrix reduces to that in the absence of Rabi coupling, independent of the parameters of the Rabi drive, with $T_{-}(\epsilon_{-}+\epsilon_{\bf k}+\epsilon_{{\bf k}3})=\frac{2\pi}{m_{r}}\frac{1}{a^{-1}+ik}$.
@@
Figure 3(a) shows the dressed scattering length and associated resonances, calculated for three sets of bare scattering lengths that are located in each colored area of the existence diagram, as indicated by the symbols (,$\bullet$,$\blacklozenge$,$\blacksquare$ , ◆ , ■) in Fig. 1(b). We see that the dressed scattering length exhibits a single resonance when only one of the bare scattering lengths is positive, while there are two resonances when both of the bare scattering lengths are positive, as expected from the existence diagram and the discussion above. One can also understand the behavior of $a_{-}$ in the vicinity of the resonances, which we find can be approximated as
@@
where we have introduced an effective background scattering length $a_{\text{bg}}=a_{13}c^{2}+a_{23}s^{2}$, and $\epsilon_{\pm},c,s$ are evaluated at the critical detuning $\delta_{c}$. Hence, we expect the resonance width $\Gamma$ to be reduced when $|a_{13}-a_{23}|$ decreases, as can be observed in Fig. 3(a), and to be enhanced when $\Omega$ increases. Interestingly, we see that Eq. (11a) resembles the typical two-channel behavior of a Feshbach resonance [45, 46, 1], although in the Rabi-driven case, $a_{\text{bg}}$ also depends on the detuning. It also has the same form as the real part of the (inverse) effective scattering length in the presence of an oscillating magnetic field [47]. However, in our scenario the scattering length is purely real, and unlike the scenario in Ref. [47] the background scattering length itself depends on the applied field.
@@
In contrast to $a_{-}$, it does not exhibit any resonance and it is in general complex valued since the scattering takes place in the dressed continuum (except if $a_{13}=a_{23}=a$, in which case it reduces to $a_{+}=a$). When $\sqrt{2m_{r}(\epsilon_{+}-\epsilon_{-})}(a_{13}c^{2}+a_{23}s^{2})\ll 1$, the effect of the continuum is less relevant and we have $a_{+}\simeq(a_{13}s^{2}+a_{23}c^{2})$.
@@
Figure 3(b) illustrates the case where only one scattering length is positive. In this case, we find a single bound state below the continuum (shaded purple area starting at $\epsilon_{-}$). At negative detuning the bound state approaches the bare bound state energy, $-1/(2m_{r}a_{23}^{2})+\delta$, from below. On the other hand, when the hybrid bound state approaches the continuum at positive detuning, its energy is modified. The detuning at which the bound state enters the continuum corresponds to the position of a resonance in Fig. 3(a).
@@
Finally, we comment on the stability of the Rabi-coupled Fermi gas. In general, strongly interacting three-component Fermi gases are unstable towards three-body recombination [30, 12]. While this is not an issue for the two-body physics described above, as it only involves two particles, the Rabi-dressed many-body system necessarily features all three bare fermionic components. However, we note that, in our setup, the absence of 1-2 interactions strongly suppresses both Efimov physics and three-body losses compared with the case where all components interact near resonantly [57]. Additional stability can be achieved by considering $\Omega\gtrsim E_{F},T$ such that the population of the upper dressed state remains small and we are left with an effective two-component Fermi gas. Indeed, the polaron limit of the strongly Rabi-driven Fermi gas was investigated in Ref. [29], and despite the significant impurity fraction of 1:5 the authors did not find that losses noticeably impacted their measurements.
@@ Appendix A SUPPLEMENTAL MATERIAL: Scattering resonances and pairing in a Rabi-coupled Fermi gas
A.1 Details on the calculation of the $T$ matrix
@@
$\Pi_{\sigma\sigma^{\prime}}(\omega)$ $=\int\frac{d\omega^{\prime}}{2i\pi}\sum_{\bf q}G_{\sigma\sigma^{\prime}}(-{\bf q},-\omega^{\prime})G_{3}({\bf q},\omega+\omega^{\prime}).$ (S2)
The non-interacting Green’s function of the third species is given by $G_{3}({\bf k},\omega)=(\omega-\epsilon_{{\bf k}3}+i0)^{-1}$, and those of the Rabi-coupled species $G_{\sigma\sigma^{\prime}}({\bf k},\omega)=\bra{0}\hat{f}_{{\bf k}\sigma^{\prime}}(\omega-\hat{H}_{0}+i0)^{-1}\hat{f}_{{\bf k}\sigma}^{\dagger}\ket{0}$ read [69, 42]
@@
A.1.1 Many-body $T$ matrix
@@
A.2 Bound states using the Schrödinger equation
@@
|$\ket{\psi}$ $=\sum_{\bf k}\varphi_{\bf k}\hat{f}_{{\bf k}3}^{\dagger}\left(c_{\bf k}\hat{f}_{-{\bf k}1}^{\dagger}+s_{\bf k}\hat{f}_{-{\bf k}2}^{\dagger}\right)\ket{0}.$ (S10)
Note that $c_{\bf k}$ and $s_{\bf k}$ are part of the bound state wave function that we wish to determine and differ from the single-particle mixing coefficients. Projecting the two-body Schrödinger equation $\hat{H}\ket{\psi}=E\ket{\psi}$ onto its components, we obtain
@@
Multiplying (S11a) by $g_{13}$ and (S11b) by $g_{23}$, dividing both sides by $(E-...)$ and summing over ${\bf k}$, these equations can be rearranged as
@@
which corresponds to the equation for the poles of the dressed $T$ matrix (6).
@@
$\mathcal{S}_{x}$ $=\frac{\sum_{\bf k}\langle\hat{f}_{{\bf k}1}^{\dagger}\hat{f}_{{\bf k}2}+\hat{f}_{{\bf k}2}^{\dagger}\hat{f}_{{\bf k}1}\ket{\psi}}{\sum_{\bf k}\bra{\psi}\hat{f}_{{\bf k}1}^{\dagger}\hat{f}_{{\bf k}1}+\hat{f}_{{\bf k}2}^{\dagger}\hat{f}_{{\bf k}2}\ket{\psi}}.$ (S21b)

### v2

@@ Abstract
[…]
with $\sigma,\sigma^{\prime}=\{1,2\}$, are not diagonal and depend on the Rabi drive (the factor +i0+i0 shifts poles infinitesimally into the lower half place). The diagrams corresponding to the Rabi-dressed Green’s functions are presented in Fig. 2(a,b). By contrast, the third species is unaffected by the Rabi drive and thus its single-particle Green’s function is simply given by $G_{3}({\bf k},\omega)=(\omega-\epsilon_{{\bf k}3}+i0)^{-1}$.
@@
We now turn to the interactions between an atom of species 3 with a Rabi-coupled atom, Fig. 1(a). The Rabi-induced interconversion processes discussed above must be accounted for, since it can couple bound and scattering states. As a result, we find that the two-body $T$-matrix that describes the interactions can be conveniently written as a $2\times 2$ matrix equation of the form
@@
$\mathbf{g}$ $=\begin{pmatrix}g_{13}&0\\ 0&g_{23}\end{pmatrix},\penalty 10000\ \penalty 10000\ \mathbf{\Pi}(\omega)=\begin{pmatrix}\Pi_{11}(\omega)&\Pi_{12}(\omega)\\ \Pi_{21}(\omega)&\Pi_{22}(\omega)\end{pmatrix}$, (7)
@@
$\Pi_{\sigma\sigma^{\prime}}(\omega)$ $={\color[rgb]{0,0,0}i}\int\frac{d\omega^{\prime}}{2\pi}\sum_{\bf q}G_{\sigma\sigma^{\prime}}(-{\bf q},-\omega^{\prime})G_{3}({\bf q},\omega+\omega^{\prime}).$ (8)
@@
Our model Hamiltonian (1) does not include interactions between species 1 and 2. This is motivated by the fact that these are irrelevant for the two-body physics we discuss here and below, since it involves a single atom in a superposition of the Rabi-driven 1-2 components. Furthermore, we expect that a small $a_{12}$ will not qualitatively affect our results regarding the pairing instabilities at finite density as long as the system remains in the strongly driven regime $\Omega\gtrsim E_{F}$, where the pairing between species 1 and 2 can be neglected.
Figure 3: Rabi-dressed Feshbach resonances and hybrid bound states as a function of detuning $\delta/\Omega$. (a) Dressed scattering length for three sets of bare scattering lengths ($a_{13},a_{23}$) corresponding to the points marked by ($\bullet,\blacklozenge,\blacksquare$) symbols in Fig. 1(b). (b,c) Energy spectrum for $(a_{13},a_{23})\sqrt{2m_{r}\Omega}=(-0.3,0.8)$ in (b), and $(a_{13},a_{23})\sqrt{2m_{r}\Omega}=(0.6,0.8)$ in (c). We show the hybrid bound states (black lines), with the dashed red and blue lines corresponding to the bare 1-3 and 2-3 bound states. The purple lines show the dressed single-particle energies $\epsilon_{\pm}$ and the shaded area highlights the scattering continuum of the driven system.
@@
In particular, we see that in general $a_{-}\neq a_{13}c^{2}+a_{23}s^{2}$, and that the effective range $r_{\text{eff}}\leq 0$, similar to the case for a standard two-channel model of a magnetically tunable Feshbach resonance [40, 41]. We also note that when $a_{23}=0$, Eq. (10a) reduces to the effective scattering length derived in Refs. [42, 43]. In the high-symmetry scenario where $a_{13}=a_{23}=a$, Eq. (10a) reduces to $a_{-}=a$. This can be understood from the fact that, in this limit, the parts of the Hamiltonian corresponding to the lower and upper dressed fermions decouple [38]. Thus for this special case, the dressed scattering length is independent of the Rabi drive and only diverges when $a$ is itself divergent  [^1] Indeed, when $a_{13}=a_{23}=a$, the $T$ matrix reduces to that in the absence of Rabi coupling, independent of the parameters of the Rabi drive, with $T_{-}(\epsilon_{-}+\epsilon_{\bf k}+\epsilon_{{\bf k}3})=\frac{2\pi}{m_{r}}\frac{1}{a^{-1}+ik}$.
@@
Figure 3(a) shows the dressed scattering length and associated resonances, calculated for three sets of bare scattering lengths that are located in each colored area of the existence diagram, as indicated by the symbols ($\bullet,\blacklozenge,\blacksquare$) in Fig. 1(b). We see that the dressed scattering length exhibits a single resonance when only one of the bare scattering lengths is positive, while there are two resonances when both of the bare scattering lengths are positive, as expected from the existence diagram and the discussion above. One can also understand the behavior of $a_{-}$ in the vicinity of the resonances, which we find can be approximated as
@@
where we have introduced an effective background scattering length $a_{\text{bg}}=a_{13}c^{2}+a_{23}s^{2}$, and $\epsilon_{\pm},c,s$ are evaluated at the critical detuning $\delta_{c}$. Hence, we expect the resonance width ${\color[rgb]{0,0,0}|}\Gamma{\color[rgb]{0,0,0}|}$ to be reduced when $|a_{13}-a_{23}|$ decreases, as can be observed in Fig. 3(a), and to be enhanced when $\Omega$ increases. We note that $\Gamma$ can be either positive or negative, and that both cases appear in Fig. 3(a) where we can observe inverted resonance shapes on the negative and positive detuning sides. Interestingly, we see that Eq. (11a) resembles the typical two-channel behavior of a Feshbach resonance [45, 46, 1], although in the Rabi-driven case, $a_{\text{bg}}$ also depends on the detuning. It also has the same form as the real part of the (inverse) effective scattering length in the presence of an oscillating magnetic field [47]. However, in our scenario the scattering length is purely real, and unlike the scenario in Ref. [47] the background scattering length itself depends on the applied field.
@@
In contrast to $a_{-}$, it does not exhibit any resonance and it is in general complex valued since the scattering takes place in the dressed continuum (except if $a_{13}=a_{23}=a$, in which case it reduces to $a_{+}=a$). The effect of the continuum is less relevant when $\sqrt{2m_{r}(\epsilon_{+}-\epsilon_{-})}(a_{13}c^{2}+a_{23}s^{2})\ll 1$ where we have $a_{+}\simeq(a_{13}s^{2}+a_{23}c^{2})$, or in the limits $\delta/\Omega\rightarrow\pm\infty$ where $a_{+}$ reduces to $a_{23}$ or $a_{13}$.
@@
Figure 3(b) illustrates the case where only one scattering length is positive. In this case, we find a single bound state below the continuum (shaded purple area starting at $\epsilon_{-}$). At negative detuning the bound state approaches the bare bound state energy, $1/(2m_{r}a_{23}^{2})+\delta$, from below. On the other hand, when the hybrid bound state approaches the continuum at positive detuning, its energy is modified. The detuning at which the bound state enters the continuum corresponds to the position of a resonance in Fig. 3(a).
@@
Finally, we comment on the stability of the Rabi-coupled Fermi gas. In general, strongly interacting three-component Fermi gases are unstable towards three-body recombination [30, 12]. While this is not an issue for the two-body physics described above, as it only involves two particles, the Rabi-dressed many-body system necessarily features all three bare fermionic components. However, we note that, in our setup, the negligible 1-2 interactions strongly suppresses both Efimov physics and three-body losses compared with the case where all components interact near resonantly [57]. Additional stability can be achieved by considering $\Omega\gtrsim E_{F},T$ such that the population of the upper dressed state remains small and we are left with an effective two-component Fermi gas. Indeed, the polaron limit of the strongly Rabi-driven Fermi gas was investigated in Ref. [29], and despite the significant impurity fraction of 1:5 the authors did not find that losses noticeably impacted their measurements.
@@ Acknowledgements.
Data Availability.— The data that support the findings of this article are openly available [68].
@@ Appendix A SUPPLEMENTAL MATERIAL: Scattering resonances and pairing in a Rabi-coupled Fermi gas
A.1 Absence of Rabi-induced resonances when $a_{13}=a_{23}$
In this section, we show that the absence of resonances when the bare scattering lengths are equal can be understood from the symmetry of the Hamiltonian. To visualize this, it is convenient to write the interaction part of the Hamiltonian in terms of the dressed fermions operators:
$\hat{V}$ $=\hat{V}_{1}+\hat{V}_{2}$
$=g_{13}\sum_{{\bf k},{\bf k}^{\prime},{\bf q}}\hat{f}_{{\bf k}+{\bf q}1}^{\dagger}\hat{f}_{{\bf k}1}\hat{f}_{{\bf k}^{\prime}-{\bf q}3}^{\dagger}\hat{f}_{{\bf k}^{\prime}3}+g_{23}\sum_{{\bf k},{\bf k}^{\prime},{\bf q}}\hat{f}_{{\bf k}+{\bf q}2}^{\dagger}\hat{f}_{{\bf k}2}\hat{f}_{{\bf k}^{\prime}-{\bf q}3}^{\dagger}\hat{f}_{{\bf k}^{\prime}3}$ (S1a)
$=(g_{13}c^{2}+g_{23}s^{2})\sum_{{\bf k},{\bf k}^{\prime},{\bf q}}\hat{f}_{{\bf k}+{\bf q}-}^{\dagger}\hat{f}_{{\bf k}-}\hat{f}_{{\bf k}^{\prime}-{\bf q}3}^{\dagger}\hat{f}_{{\bf k}^{\prime}3}+(g_{13}s^{2}+g_{23}c^{2})\sum_{{\bf k},{\bf k}^{\prime},{\bf q}}\hat{f}_{{\bf k}+{\bf q}+}^{\dagger}\hat{f}_{{\bf k}+}\hat{f}_{{\bf k}^{\prime}-{\bf q}3}^{\dagger}\hat{f}_{{\bf k}^{\prime}3}$
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ +cs(g_{13}-g_{23})\sum_{{\bf k},{\bf k}^{\prime},{\bf q}}\left(\hat{f}_{{\bf k}+{\bf q}+}^{\dagger}\hat{f}_{{\bf k}-}+\hat{f}_{{\bf k}+{\bf q}-}^{\dagger}\hat{f}_{{\bf k}+}\right)\hat{f}_{{\bf k}^{\prime}-{\bf q}3}^{\dagger}\hat{f}_{{\bf k}^{\prime}3},$ (S1b)
where we have used the transformation given in Eq. (4) of the main text. As we can see from the second line in Eq. (S1b), in the dressed basis, the interactions with the third species give rise to an effective coupling between the dressed fermions $\hat{f}_{\pm}$. However, this coupling vanishes when $g_{13}=g_{23}=g$, and thus, in this special case, the three-component system can be effectively reduced to two two-component systems $\{-,3\}$ and $\{+,3\}$ with interaction strengths $g$. Since the $\{-,3\}$ and $\{+,3\}$ scattering channels are now completely decoupled, the presence of a bound state in the $\{+,3\}$ channel cannot affect the scattering in the $\{-,3\}$ channel, and thus, it cannot give rise to new resonances. This is also why the dressed $T$-matrices reduce to $T_{\pm}(\omega)=\frac{2\pi/m_{r}}{a^{-1}-\sqrt{\epsilon_{\pm}-\omega-i0}}$, i.e., they take the same form as a standard bare $T$-matrix with a shifted continuum.
A.2 Details on the calculation of the $T$ matrix
@@
$\Pi_{\sigma\sigma^{\prime}}(\omega)$ $={\color[rgb]{0,0,0}i}\int\frac{d\omega^{\prime}}{2\pi}\sum_{\bf q}G_{\sigma\sigma^{\prime}}(-{\bf q},-\omega^{\prime})G_{3}({\bf q},\omega+\omega^{\prime}).$ (S3)
[…]
@@
A.2.1 Many-body $T$ matrix
@@
A.3 Bound states using the Schrödinger equation
@@
[…]
@@
Multiplying (S12a) by $g_{13}$ and (S12b) by $g_{23}$, dividing both sides by (E)(E-...) and summing over $\bf k$}, these equations can be rearranged as
@@
which corresponds to the equation for the poles of the dressed $T$ matrix (S2).
@@
[…]